Write the Golden Ratio up to 30 decimals places.
1
Expert's answer
2020-09-28T17:12:58-0400

Two quantities a and b are said to be in the golden ratio φ if

"\frac{a+b}{a}=\frac{a}{b}=\varphi"

One method for finding the value of φ is to start with the left fraction. Through simplifying the fraction and substituting in"\frac{b}{a}=\frac{1}{\varphi},"

"{\frac {a+b}{a}}={\frac {a}{a}}+{\frac {b}{a}}=1+{\frac {b}{a}}=1+{\frac {1}{\varphi }}" b/a = 1/φ,


"1+{\frac {1}{\varphi }}=\varphi"

"\varphi +1=\varphi ^{2}"

"{\varphi }^{2}-\varphi -1=0."

Using the quadratic formula, two solutions are obtained:

"x={\frac {-b\pm {\sqrt {b^{2}-4ac}}}{2a}}"


"{\frac {1+{\sqrt {5}}}{2}}=1.618\,033\,988\,749\,894\,848\,204\,586\,834\,365\dots"

and

"{\frac {1-{\sqrt {5}}}{2}}=-0.618\,033\,988\,7\dots"


Because φ is the ratio between positive quantities, φ is necessarily positive:

"\varphi ={\frac {1+{\sqrt {5}}}{2}}=1.618\,033\,988\,749\,894\,848\,204\,586\,834\,365\dots"